OpenStudy (anonymous):

F''''(x) = sinx F'''(-pi/2)=0 F''(0)=0 F'(pi/2)=0 F(-2012pi)=2012 Find F(x):

OpenStudy (anonymous):

keep integrating

OpenStudy (roseflight):

if F''''=sin(x) then F''' would be -cos(x)+c. Since you have F'''(-pi/2)=0 though you need to solve for c: 0=-cos(-pi/2)+c c=0 good, we can now say that F'''=-cos(x) do the same thing with F'' then (F''=-sin(x)+c, 0=-sin(0)+c, c=0, etc.)